The sum of two numbers is 26. Their difference is 4. Find the smaller number.
Found 2 solutions by stanbon, lynnlo:
Answer by stanbon(75887)

The sum of two numbers is 26. Their difference is 4. Find the smaller number.
----
Equations:
x + y = 26
x - y = 4
---
Add to get:
2x = 30
x = 15
---
Solve for "y":
x + y = 26
15 + y = 26
y = 11 (smaller number)
========================
cheers,
Stan H.

Answer by lynnlo(4176)

the smaller number is 11